mallll:

Factor the GCF: 21x3y4 + 15x2y2 − 12xy3. 3xy(7x2y3 + 5x − 4y) 3xy(7x2y + 5x − 4y2) 3xy2(7x2y + 5x − 4y) 3xy2(7x2y2 + 5x − 4y)

jhonyy9:

hi any idea ? do you know what mean GCF ?

jhonyy9:

exactly - nice so from these choices what you think that may be true ?

jhonyy9:

ok step by step for 21, 15 and 12 what nr. may be the GCF ?

jhonyy9:

are you sure ?

jhonyy9:

why not ? bc. this is right

mallll:

ok i thought it was wrong because i guesses to be honest :/

jhonyy9:

so but do you know why is right ? i will try explain it to you - just please collaborate - ok?

mallll:

ok

jhonyy9:

ty. the GCF for 12,15 and 21 what nr. is ?

jhonyy9:

how you can get it ?

jhonyy9:

do you know the way ?

mallll:

not really but im pretty sure its 3

jhonyy9:

ok looke please 21 = 3*7 15 = 3*5 12 = 3*4 so what nr. is common there ?

mallll:

3

jhonyy9:

yes so please remember and try learn it tis is the way ok. ? now for x^3 ,x^2 and x what is the common term ?

jhonyy9:

x^3 = x*x*x x^2 =x*x x = x

mallll:

so the term would be x

jhonyy9:

yes sure and now for y there are y^4 = y^2 *y^2 y^2 =y^2 y^3 = y*y^2

mallll:

so that term would be y right

jhonyy9:

yes y is common term but you need getting the greatest

mallll:

so ho would i do that ??

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 yes sure and now for y there are y^4 = y^2 *y^2 y^2 =y^2 y^3 = y*y^2 \(\color{#0cbb34}{\text{End of Quote}}\) above than you check what term of y is common ?

mallll:

2

jhonyy9:

no y^2 not is greatest than y ?